mr alvarado is 20 years older than his son . in 5 years, he will be twice as old as his son will then be find their present ags
Answer by richwmiller(17219)

d=s+20,
d+5=2*(s+5)
s+20+5=2s+10
s+25=2s+10
15=s
35=d